SYZYGIES OF UNIMODULAR LAWRENCE IDEALS

DAVE BAYER SORIN POPESCU BERND STURMFELS

Abstract. Infinite hyperplane arrangements whose vertices form a lattice are studied from the point of view of commutative algebra. The quotient of such an arrangement modulo the lattice action represents the minimal free resolution of the associated binomial ideal, which defines a toric subvariety in a product of projective lines. Connections to graphic arrangements and to Beilinson’s spectral sequence are explored.